If the input impedance is defined as Zin = Vs/s, find the input impedance of the op amp circuit given below when R1 = 20 kN,Asap please, I willl upvote
Answer #1

JX5oooxlox 109 Sol - Xcia Xol If IWC Vo - 20000 vo Rp Up R22 Xc2=1 Is Ous guce = Jxsöport Jx5800x20x109 Xca Xc2=-10000] Zin pfrom egn putting to value in ear Vp (2+J) VP ( 2+1) = Vs + Vo (19) Vs + VP (145) 1+27 Vp (2+5 -(0.6 -0.2J)) = VS Vel 2+-0.6 +